OpenStudy (kohai):

The height of a grain of a cylindrical silo is is increasing at a constant rate of 4 feet per minute At what rate is the volume of grain in the cylinder if the radius of the silo is 10 feet?

ganeshie8 (ganeshie8):

amount of grain accumulated in the cylinder at a given height :\[\large \rm v = \pi r^2 h = 100\pi h \]

ganeshie8 (ganeshie8):

differentiate both sides with respect to time

OpenStudy (kohai):

I know the process theoretically, it's just.. doing it.. that I have issues with..

ganeshie8 (ganeshie8):

that means you need to do it more and more till you feel at peace with the process

ganeshie8 (ganeshie8):

Notice that volume depends on height and the height depends on time

ganeshie8 (ganeshie8):

the question gave you how the height of grains poured into the container is depending on time

ganeshie8 (ganeshie8):

you need to find out how volume depends on time

OpenStudy (kohai):

So dV/dt essentially?

ganeshie8 (ganeshie8):

Exactly!

OpenStudy (kohai):

Ok. I'm going to work on it and check back with you :)

ganeshie8 (ganeshie8):

k

OpenStudy (kohai):

Okay

OpenStudy (kohai):